# Gaussian Mixture Models Based Augmentation Enhances GNN Generalization

### v2

@@ Abstract
Graph Neural Networks (GNNs) have shown great promise in tasks like node and graph classification, but they often struggle to generalize, particularly to unseen or out-of-distribution (OOD) data. These challenges are exacerbated when training data is limited in size or diversity. To address these issues, we introduce a theoretical framework using Rademacher complexity to compute a regret bound on the generalization error and then characterize the effect of data augmentation. This framework informs the design of GMM-GDA, an efficient graph data augmentation (GDA) algorithm leveraging the capability of Gaussian Mixture Models (GMMs) to approximate any distribution. Our approach not only outperforms existing augmentation techniques in terms of generalization but also offers improved time complexity, making it highly suitable for real-world applications. Our code is publicly available at: https://github.com/abbahaddou/GMM-GDA.
@@ 1 Introduction
Graphs are a fundamental and ubiquitous structure for modeling complex relationships and interactions. In biology, graphs are employed to represent complex networks of protein interactions and in drug discovery by modeling molecular relationships. Similarly, in social networks, graphs capture relationships and community interactions, offering insights into social structures and interactions (Zeng et al., 2022; Gaudelet et al., 2021; Newman et al., 2002). To address the unique challenges posed by graph-structured data, GNNs have been developed as a specialized class of neural networks designed to operate directly on graphs. Unlike traditional neural networks that are optimized for grid-like data, such as images or sequences, GNNs are engineered to process and learn from the relational information embedded in graph structures. GNNs have demonstrated state-of-the-art performance across a range of graph representation learning tasks such as node and graph classification, proving their effectiveness in various real-world applications (Vignac et al., 2022; Corso et al., 2022; Duval et al., 2023; Castro-Correa et al., 2024; Chi et al., 2022).
@@
Despite their impressive capabilities, GNNs face significant challenges related to generalization, particularly when handling unseen or out-of-distribution (OOD) data (Guo et al., 2024; Li et al., 2022). OOD graphs are those that differ significantly from the training data in terms of graph structure, node features, or edge types, making it difficult for GNNs to adapt and perform well on such data. This challenge is also faced when GNNs are trained on small datasets, where the limited data diversity hampers the model’s ability to generalize effectively. To address these challenges, the community has explored various strategies to improve the robustness and generalization ability of GNNs (Abbahaddou et al., 2024; Yang et al., 2022). One promising approach is data augmentation, which involves artificially expanding the training dataset by introducing variations of the original graph data. Data augmentation has shown its benefits across different types of data structures such as images (Krizhevsky et al., 2012) and time series (Aboussalah et al., 2023). For graph data structures, generating augmented versions of the original graphs, such as by adding or removing nodes and edges or perturbing node features (Rong et al., 2019; You et al., 2020), allows for the creation of a more varied training set. Inspired by the success of the Mixup technique in computer vision (Rebuffi et al., 2021; Dabouei et al., 2021; Hong et al., 2021), additional methods such as $\mathcal{G}$-Mixup and GEOMIX have been developed to adapt the Mixup technique for graph data (Ling et al., 2023; Han et al., 2022). These techniques combine different graphs to create new, synthetic training examples, further enriching the dataset and enhancing the GNN’s ability to generalize to new unseen graph structures.
@@
In this work, we introduce a novel graph augmentation technique based on Gaussian Mixture Models (GMMs), which operates at the level of the final hidden representations. Specifically, guided by our theoretical results, we apply the Expectation-Maximization (EM) algorithm to train a GMM on the graph representations. We then use this GMM to generate new augmented graph representations through sampling, enhancing the diversity of the training data.
Contributions. The contributions of our work are as follows:
@@
  • Theoretical framework for generalization in GNNs: We introduce a theoretical framework that rigorously analyzes how graph data augmentation impacts the generalization capabilities of GNNs. This framework offers new insights into the underlying mechanisms that drive performance improvements through augmentation.
@@
  • Efficient graph data augmentation via GMMs: We propose GMM-GDA, a fast and efficient graph data augmentation technique, leveraging GMMs. This approach enhances the diversity of training data while maintaining computational simplicity, making it scalable for large graph datasets.
@@
  • Comprehensive theoretical analysis using influence functions: We perform an in-depth theoretical analysis of our augmentation strategy through the lens of influence functions, providing a principled understanding of the approach’s impact on generalization performance.
@@ 2 Background and Related Work
Graph Neural Networks. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph, where $\mathcal{V}$ represents the set of vertices and $\mathcal{E}$ represents the set of edges. We use $L=|\mathcal{V}|$ to denote the number of vertices and $m=|\mathcal{E}|$ to denote the number of edges. For a node $v\in V$, let $\mathcal{N}(v)$ be the set of its neighbors, defined as $\mathcal{N}(v)=\{u\colon(v,u)\in E\}$. The degree of vertex $v$ is the number of neighbors it has, which is $|\mathcal{N}(v)|$. A graph is commonly represented by its adjacency matrix $\mathbf{A}\in\mathbb{R}^{L\times L}$, where the $(i,j)$-th element of this matrix is equal to the weight of the edge between the $i$-th and $j$-th node of the graph and a weight of zero in case the edge does not exist. Additionally, in some cases, nodes may have associated feature vectors. We denote these node features by $\mathbf{X}\in\mathbb{R}^{L\times D}$ where $D$ is the dimensionality of the features.
@@
A GNN model consists of multiple neighborhood aggregation layers that use the graph structure and the feature vectors from the previous layer to generate updated representations for the nodes. Specifically, GNNs update a node’s feature vector by aggregating information from its local neighborhood. Consider a GNN model with $T$ neighborhood aggregation layers. Let $\mathbf{h}_{v}^{(0)}$ denote the initial feature vector of node $v$, which is the corresponding row in $\mathbf{X}$. At each layer $t>0$, the hidden state $\mathbf{h}_{v}^{(t)}$ of node $v$ is updated as follows:
$\mathbf{a}_{v}^{(t)}=\texttt{AGGREGATE}^{(t)}\Big{(}\big{\{}\mathbf{h}_{u}^{(t-1)}\colon u\in\mathcal{N}(v)\big{\}}\Big{)},$
$\mathbf{h}_{v}^{(t)}=\texttt{COMBINE}^{(t)}\Big{(}\mathbf{h}_{v}^{(t-1)},\mathbf{a}_{v}^{(t)}\Big{)},$
@@
where $\texttt{AGGREGATE}(\cdot)$ is a permutation-invariant function that combines the feature vectors of $v$’s neighbors into an aggregated vector. This aggregated vector, together with the previous feature vector $\mathbf{h}_{v}^{(t-1)}$, is fed to the $\texttt{COMBINE}(\cdot)$ function, which merges these two vectors to produce the updated feature vector of $v$. Two popular GNN architechtures are Graph Convolution Networks (GCN) and Graph Isomorphism Networks (GIN) (Kipf & Welling, 2017; Xu et al., 2019).
After $T$ iterations of neighborhood aggregation, to produce a graph-level representation, GNNs apply a permutation invariant readout function, e.g., the sum operator, to nodes feature as follows:
$\mathbf{h}_{\mathcal{G}}=\texttt{READOUT}\Big{(}\big{\{}\mathbf{h}_{v}^{(T)}\colon v\in V\big{\}}\Big{)}.$ (1)
@@
Data Augmentation for Graphs. Graph data augmentation has become an essential aspect of enhancing the performance and robustness of GNNs. Among the classical techniques, structural modifications of the graph are widely used to generate augmented training graphs. Key methods in this category include DropEdge, DropNode, and Subgraph sampling techniques (Rong et al., 2019; You et al., 2020). For example, the DropEdge technique randomly removes a subset of edges from the graph during training, improving the model’s robustness to missing or noisy connections. Similarly, DropNode removes certain nodes as well as their connections, assuming that the missing part of nodes will not affect the semantic meaning, i.e., the structural and relational information of the original graph. Another method is Subgraph, which samples a subgraph from the original graph using random walk to use as a training graph. By training on these augmented graphs, GNNs can generalize to unseen graph structures more efficiently.
@@
Beyond classical methods, recent advancements have explored more sophisticated augmentation techniques, focusing on manipulating graph embeddings and leveraging geometric properties of graphs. Following the effectiveness of the Mixup technique in computer vision (Rebuffi et al., 2021; Dabouei et al., 2021; Hong et al., 2021), several works describe variations of the Mixup for graphs. For example, the Manifold-Mixup model conducts a Mixup operation for graph classification in the embedding space. This technique interpolates between graph-level embeddings after the READOUT function, blending different graphs in the embedding space (Wang et al., 2021). $\mathcal{G}$-Mixup (Han et al., 2022) uses graphons to model the topological structures of each graph class and then interpolates the graphons of different classes, subsequently generating synthetic graphs by sampling from mixed graphons across different classes. It is important to note that $\mathcal{G}$-Mixup operates under a significant assumption: graphs belonging to the same class can be produced by a single graphon. The $S$-Mixup method, for a given pair of graphs, determines node-level correspondences between the nodes in both graphs and subsequently interpolates the graphs (Ling et al., 2023). FGW-Mixup adopts the Fused Gromov–Wasserstein barycenter as the mixup graphs, but suffers from heavy computation (Ma et al., 2024). Finally, the GeoMix technique (Zeng et al., 2024) uses mixup graphs on the exact Gromov-Wasserstein geodesics.
Gaussian Mixture Models. GMMs are probabilistic models used for modeling complex data by representing them as a mixture of multiple Gaussian distributions. The probability density function $p(\mathbf{x})$ of a data point $\mathbf{x}$ in a GMM with $K$ Gaussian components is given by:
$p(\mathbf{x})=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\mathbf{x}\mid\bm{\mu}_{k},\bm{\Sigma}_{k}),$ (2)
where $\pi_{k}$ is the weight of the $k$-th Gaussian component, with $\pi_{k}\geq 0$ and $\sum_{k=1}^{K}\pi_{k}=1$, and $\mathcal{N}(\mathbf{x}\mid\bm{\mu}_{k},\bm{\Sigma}_{k})$ is the Gaussian probability density function for the $k$-th component, defined as:
$\mathcal{N}(\mathbf{x}\mid\bm{\mu}_{k},\bm{\Sigma}_{k})=\frac{1}{(2\pi)^{d/2}\det(\bm{\Sigma}_{k})^{1/2}}\exp\left(-\frac{1}{2}(\mathbf{x}-\bm{\mu}_{k})^{\top}\bm{\Sigma}_{k}^{-1}(\mathbf{x}-\bm{\mu}_{k})\right),$
where $\bm{\mu}_{k}$ and $\bm{\Sigma}_{k}$ are respectively the mean vectors and the covariance vectors of the $k$-th Gaussian component, and $d$ the dimensionality of $\mathbf{x}$. The parameters of a GMM are typically estimated using the EM algorithm (Dempster et al., 1977), which alternates between estimating the membership probabilities of data points for each Gaussian component (Expectation step) and updating the parameters of the Gaussian distributions (Maximization step). GMMs are a powerful tool in statistics and machine learning and are used for various purposes, including clustering and density estimation (Ozertem & Erdogmus, 2011; Naim & Gildea, 2012; Zhang et al., 2021).
3 GMM-GDA: Gaussian Mixture Model for Graph Data Augmentation
@@
In this section, we begin by introducing the mathematical framework for graph data augmentation and its connection to the generalization of GNNs. Following that, we present our proposed model GMM-GDA, which is based on GMMs for graph augmentation.
3.1 Mathematical Formalism
@@
We focus on the task of graph classification, where the objective is to classify graphs into predefined categories. Given a training dataset of graphs $\mathcal{D}_{\text{train}}=\{(\mathcal{G}_{n},y_{n})\mid n=1,\ldots,N\}$, $\mathcal{G}_{n}$ is the $n$-th graph and $y_{n}$ is its corresponding label belonging to a set $\{0,\ldots,C\}$. Each graph $\mathcal{G}_{n}$ is represented as a tuple $(\mathcal{V}_{n},\mathcal{E}_{n},\mathbf{X}_{n})$, where $\mathcal{V}_{n}$ denotes the set of nodes with cardinality $L_{n}=|\mathcal{V}_{n}|$, $\mathcal{E}_{n}\subseteq\mathcal{V}_{n}\times\mathcal{V}_{n}$ is the set of edges, and $\mathbf{X}_{n}\in\mathbb{R}^{L_{n}\times D}$ is the node feature matrix of dimension $D$. The objective is to train a GNN $f(\cdot,\theta)$ that can accurately predict the class labels for unseen graphs in the test set $\mathcal{D}_{\text{test}}=\{\mathcal{G}_{n}^{test}\mid n=1,\ldots,N_{\text{test}}\}$. The classical training approach involves minimizing the following loss function,
$\mathcal{L}=\frac{1}{N}\sum_{n=1}^{N}\ell(f(\mathcal{G}_{n},\theta),y_{n}),$ (3)
@@
where $\ell$ denotes the cross-entropy loss function. To improve the robustness and generalization ability of the GNN, we introduce data augmentation for graphs. For each graph $\mathcal{G}_{n}$, we generate an augmented graph $\mathcal{G}_{n}^{\lambda}$ using a data augmentation strategy $A_{\lambda}$, where $\lambda$ is a parameter sampled from a prior distribution $\mathcal{P}$, such as a uniform distribution. The data augmentation strategy $A_{\lambda}$ is defined as a function: $A_{\lambda}:\mathcal{G}_{n}\in G\rightarrow\mathcal{G}_{n}^{\lambda}=A(\mathcal{G}_{n},\lambda)\in G$, where $G$ is the set of all possible graphs with $n$ nodes. With the augmented data, the loss function is modified to account for multiple augmented versions of each graph:
@@
$\mathcal{L}^{\text{aug}}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(f(\mathcal{G}_{n}^{\lambda},\theta),y_{n})\right].$ (4)
@@
For simplicity, we denote the loss for the original graph as $\ell(f(\mathcal{G}_{n},\theta),y_{n})=\ell(\mathcal{G}_{n},\theta)$ and the loss for an augmented graph as $\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(f(\mathcal{G}_{n}^{\lambda},\theta),y_{n})\right]=\ell_{aug}(\mathcal{G}_{n},\theta)$. The loss $\ell_{aug}$ is empirically estimated as follows,
@@
$\ell_{aug}(\mathcal{G}_{n},\theta)=\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(f(\mathcal{G}_{n}^{\lambda},\theta),y_{n})\right]\simeq\frac{1}{M}\sum_{m=1}^{M}\ell(f(\mathcal{G}_{n}^{\lambda_{n,m}},\theta),y_{n}),$ (5)
@@
where $M$ denotes the number of augmented samples per graph and $\{\lambda_{n,m}\}_{m=1}^{M}$ are the parameters sampled from $\mathcal{P}$ for each training graph $\mathcal{G}_{n}$. To understand the impact of data augmentation on the graph classification performance, we analyze the effect of sampling strategy $\mathcal{P}$ on the generalization risk $\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta)\right]$. More specifically, we want to study the generalization error $\eta=\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{aug})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{\star})\right]$, where $\theta_{aug}$ and $\theta_{\star}$ are the optimal GNN parameters for the augmented and non-augmented settings,
$\theta_{\star}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta)\right],\theta_{aug}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathcal{G}\sim G}\left[\ell_{aug}(\mathcal{G},\theta)\right]=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\theta)\right],$
and which can be estimated empirically as follows,
@@
$\hat{\theta}_{aug}$ $=\operatorname*{arg\,min}_{\theta}\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\theta)\right]\simeq\operatorname*{arg\,min}_{\theta}\frac{1}{N\times M}\sum_{n=1}^{N}\sum_{m=1}^{M}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta).$
By theoretically studying the generalization error $\eta$, we aim to quantify the effect of each augmentation strategy on the overall classification performance, providing insights into the benefits and potential trade-offs of data augmentation in graph-based learning tasks. In Theorem 3.1, we present a regret bound of the generalization error using Rademacher complexity defined as follows Yin et al. (2019),
@@
where $\epsilon_{n}$ are independent Rademacher variables, taking values $+1$ or $-1$ with equal probability, $P_{\epsilon}$ is the Rademacher distribution,and $\Theta$ is the hypothesis class. Rademacher complexity is a fundamental concept in statistical learning which indicates how well a learned function will perform on unseen data (Shalev-Shwartz & Ben-David, 2014). Lower Rademacher complexity indicates a better generalization.

### v3

# Graph Neural Network Generalization with Gaussian Mixture Model Based Augmentation

@@ Abstract
Graph Neural Networks (GNNs) have shown great promise in tasks like node and graph classification, but they often struggle to generalize, particularly to unseen or out-of-distribution (OOD) data. These challenges are exacerbated when training data is limited in size or diversity. To address these issues, we introduce a theoretical framework using Rademacher complexity to compute a regret bound on the generalization error and then characterize the effect of data augmentation. This framework informs the design of GRATIN, an efficient graph data augmentation algorithm leveraging the capability of Gaussian Mixture Models (GMMs) to approximate any distribution. Our approach not only outperforms existing augmentation techniques in terms of generalization but also offers improved time complexity, making it highly suitable for real-world applications. Our code is publicly available at: https://github.com/abbahaddou/GRATIN.
Machine Learning, ICML
@@ 1 Introduction
Graphs are a fundamental and ubiquitous structure for modeling complex relationships and interactions. In biology, graphs are employed to represent complex networks of protein interactions and in drug discovery by modeling molecular relationships (Gaudelet et al., 2021; Jagtap et al., 2022). Similarly, social networks capture relationships and community interactions (Aboussalah et al., 2023a; Zeng et al., 2022; Malliaros & Vazirgiannis, 2013; Newman et al., 2002). To address the unique challenges posed by graph-structured data, GNNs have been developed as a specialized class of neural networks designed to operate directly on graphs. Unlike traditional neural networks that are optimized for grid-like data, such as images or sequences, GNNs are engineered to process and learn from the relational information embedded in graph structures. GNNs have demonstrated state-of-the-art performance across a range of graph representation learning tasks such as node and graph classification, proving their effectiveness in various real-world applications (Vignac et al., 2023; Corso et al., 2023; Duval et al., 2023; Castro-Correa et al., 2024; Chi et al., 2022; Panagopoulos et al., 2024; Aboussalah & Ed-dib, 2025).
@@
Despite their impressive capabilities, GNNs face significant challenges related to generalization, particularly when handling unseen or out-of-distribution (OOD) data (Guo et al., 2024; Li et al., 2022). OOD graphs are those that differ significantly from the training data in terms of graph structure, node features, or edge types, making it difficult for GNNs to adapt and perform well on such data. This challenge is also faced when GNNs are trained on small datasets, where the limited data diversity hampers the model’s ability to generalize effectively. To address these challenges, the community has explored various strategies to improve the robustness and generalization ability of GNNs (Abbahaddou et al., 2024; Yang et al., 2023).
Generalization bounds for GNNs have been derived using various theoretical tools, such as the Vapnik-Chervonenkis (VC) dimension (Pfaff et al., 2021; Garg et al., 2020) and Rademacher complexity (Yin et al., 2019; Esser et al., 2021). Furthermore, Liao et al. (2021) were among the first to establish generalization bounds for GNNs using the PAC-Bayesian approach. Neural Tangent Kernels have also been employed to study the generalization properties of infinitely wide GNNs trained via gradient descent (Jacot et al., 2018; Du et al., 2019; Huang et al., 2024). While most existing research has focused on the node classification task, enhancing generalization in graph classification presents unique challenges. Techniques to improve generalization in graph classification can be broadly categorized into architectural and dataset-based strategies (Tang & Liu, 2023; Buffelli et al., 2022). On the dataset side, techniques like adversarial training and data augmentation play a significant role. More broadly, data augmentation methods create synthetic or modified graph instances to enrich the training set, reducing overfitting and enhancing the model’s adaptability to diverse graph structures. Data augmentation has shown its benefits across different types of data structures such as images (Krizhevsky et al., 2012) and time series (Aboussalah et al., 2023b). For graph data structures, generating augmented versions of the original graphs, such as by adding or removing nodes and edges or perturbing node features (Rong et al., 2020; You et al., 2020), allows for the creation of a more varied training set. Inspired by the success of the Mixup technique in computer vision (Rebuffi et al., 2021; Dabouei et al., 2021; Hong et al., 2021), additional methods such as $\mathcal{G}$-Mixup and GeoMix have been developed to adapt Mixup for graph data (Ling et al., 2023; Han et al., 2022). These techniques combine different graphs to create new, synthetic training examples, further enriching the dataset and enhancing the GNN’s ability to generalize to new unseen graph structures.
@@
In this work, we introduce GRATIN, a graph augmentation technique based on Gaussian Mixture Models (GMMs), which operates at the level of the final hidden representations. Specifically, guided by our theoretical results, we apply the Expectation-Maximization (EM) algorithm to train a GMM on the graph representations. We then use this GMM to generate new augmented graph representations through sampling, enhancing the diversity of the training data.
The contributions of our work are as follows:
@@
  • Theoretical framework for generalization in GNNs. We introduce a theoretical framework that allows us to rigorously analyze how graph data augmentation impacts the generalization GNNs. This framework offers new insights into the underlying mechanisms that drive performance improvements through augmentation.
@@
  • Efficient graph data augmentation via GMMs. We propose GRATIN, a fast and efficient graph data augmentation technique leveraging GMMs. This approach enhances the diversity of training data while maintaining computational simplicity, making it scalable for large graph datasets.
@@
  • Comprehensive theoretical analysis using influence functions. We perform an in-depth theoretical analysis of our augmentation strategy through the lens of influence functions, providing a principled understanding of the approach’s impact on generalization performance.
  • Empirical Validation. Through experiments on real-world datasets we confirm GRATIN to be a fast, high-performing graph augmentation scheme in practice.
@@ 2 Background and Related Work
Notation. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph, where $\mathcal{V}$ represents the set of vertices and $\mathcal{E}$ represents the set of edges. We use $p=|\mathcal{V}|$ to denote the number of nodes. For a node $v\in V$, let $\mathcal{N}(v)$ be the set of its neighbors, defined as $\mathcal{N}(v)=\{u\colon(v,u)\in\mathcal{E}\}$. The degree of vertex $v$ is the number of neighbors it has, which is $\text{{deg}}(v)=|\mathcal{N}(v)|$. A graph is commonly represented by its adjacency matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$, where the $(i,j)$-th element of this matrix is equal to the weight of the edge between the $i$-th and $j$-th node of the graph and a weight of zero in case the edge does not exist. Additionally, in some cases, nodes may have associated feature vectors. We denote these node features by $\mathbf{X}\in\mathbb{R}^{p\times d}$ where $d$ is the dimensionality of the features.
@@
Graph Neural Networks (GNNs). A GNN model consists of multiple neighborhood aggregation layers that use the graph structure and the feature vectors from the previous layer to generate updated representations for the nodes. Specifically, GNNs update a node’s feature vector by aggregating information from its local neighborhood. Consider a GNN model with $T$ neighborhood aggregation layers. Let $\mathbf{h}_{v}^{(0)}$ denote the initial feature vector of node $v$, which is the corresponding row in $\mathbf{X}$. At each layer $t>0$, the hidden state $\mathbf{h}_{v}^{(t)}$ of node $v$ is updated as follows:
$\mathbf{m}_{v}^{(t)}=\texttt{AGGREGATE}^{(t)}\Big{(}\big{\{}\mathbf{h}_{u}^{(t-1)}\colon u\in\mathcal{N}(v)\big{\}}\Big{)},$
$\mathbf{h}_{v}^{(t)}=\texttt{COMBINE}^{(t)}\Big{(}\mathbf{h}_{v}^{(t-1)},\mathbf{m}_{v}^{(t)}\Big{)},$
@@
where $\texttt{AGGREGATE}(\cdot)$ is a permutation-invariant function that combines the feature vectors of $v$’s neighbors into an aggregated vector. This aggregated vector, together with the previous feature vector $\mathbf{h}_{v}^{(t-1)}$, is fed to the $\texttt{COMBINE}(\cdot)$ function, which merges these two vectors to produce the updated feature vector of $v$.
After $T$ iterations of neighborhood aggregation, a GNN typically produces a graph-level representation by first applying a permutation-invariant readout function, e.g., a Sum operator, to the final node embeddings. This aggregated output is subsequently passed through a trainable neural network, such as a multi-layer perceptron (MLP), denoted by $\Psi$ and referred to as the post-readout neural network, to produce the final graph-level predictions or representations. Mathematically, this process can be expressed as
$\mathbf{h}_{\mathcal{G}}=\Psi\circ\texttt{READOUT}\Big{(}\big{\{}\mathbf{h}_{v}^{(T)}\colon v\in\mathcal{V}\big{\}}\Big{)}.$
Two popular GNN architectures are Graph Convolution Networks (GCN) and Graph Isomorphism Networks (GIN) (Kipf & Welling, 2017; Xu et al., 2019). The exact expression of these models can be found in Appendix D.
@@
Data Augmentation for Graphs. Graph data augmentation has become essential to enhance the performance and robustness of GNNs. Classical graph augmentation techniques focus on structural modifications to generate augmented graphs. Key methods here include DropEdge, DropNode, and Subgraph sampling (Rong et al., 2020; You et al., 2020). For instance, DropEdge randomly removes a subset of edges from the graph during training, improving the model’s robustness to missing or noisy connections. Similarly, DropNode removes certain nodes as well as their connections, assuming that the missing part of nodes will not affect the semantic meaning, i.e., the structural and relational information of the original graph. Subgraph sampling, on the other hand, samples a subgraph from the original graph using random walks to use as a training graph.
@@
Beyond classical methods, recent advancements have explored more sophisticated augmentation techniques, focusing on manipulating graph embeddings and leveraging the geometric properties of graphs. Following the effectiveness of the Mixup technique in computer vision (Rebuffi et al., 2021; Dabouei et al., 2021; Hong et al., 2021), several works describe variations of the Mixup for graphs. For example, the Manifold-Mixup model conducts a Mixup operation for graph classification in the embedding space. This technique interpolates between graph-level embeddings after the READOUT function, blending different graphs in the embedding space (Wang et al., 2021). Similarly, $\mathcal{G}$-Mixup (Han et al., 2022) uses graphons to model the topological structures of each graph class and then interpolates the graphons of different classes, subsequently generating synthetic graphs by sampling from mixed graphons across different classes. It is important to note that $\mathcal{G}$-Mixup operates under a significant assumption: graphs belonging to the same class can be produced by a single graphon. Other advanced techniques include $S$-Mixup method, which interpolates graphs by first determining node-level correspondences between a pair of graphs (Ling et al., 2023), and FGW-Mixup, which adopts the Fused Gromov-Wasserstein barycenter to compute mixup graphs but suffers from heavy computation time (Ma et al., 2024). Finally, GeoMix (Zeng et al., 2024) leverages Gromov-Wasserstein geodesics to interpolate graphs more efficiently. By leveraging these structural augmentation techniques, GNNs can better generalize to unseen graph structures.
3 GRATIN: Gaussian Mixture Model for Graph Data Augmentation
@@
In this section, we introduce the mathematical framework for graph data augmentation and its connection to the generalization of GNNs. Then, we present our proposed model GRATIN, which is based on GMMs for graph augmentation.
3.1 Formalism of Graph Data Augmentation
@@
We focus on the task of graph classification, where the objective is to classify graphs into predefined categories. Let $\mathcal{D}$ denote the distribution of graphs. Given a training set of graphs $\mathcal{D}_{\text{train}}=\{(\mathcal{G}_{n},y_{n})\mid n=1,\ldots,N\}$, $\mathcal{G}_{n}$ is the $n$-th graph and $y_{n}$ is its corresponding label belonging to a set $\{0,\ldots,C\}$. Each graph $\mathcal{G}_{n}$ is represented as a tuple $(\mathcal{V}_{n},\mathcal{E}_{n},\mathbf{X}_{n})$, where $\mathcal{V}_{n}$ denotes the set of nodes with cardinality $p_{n}=|\mathcal{V}_{n}|$, $\mathcal{E}_{n}\subseteq\mathcal{V}_{n}\times\mathcal{V}_{n}$ is the set of edges, and $\mathbf{X}_{n}\in\mathbb{R}^{p_{n}\times d}$ is the node feature matrix of dimension $d$. The objective is to train a GNN $f(\cdot,\theta)$ that can accurately predict the class labels for unseen graphs in the test set $\mathcal{D}_{\text{test}}=\{\mathcal{G}_{n}^{\text{test}}\mid n=1,\ldots,N_{\text{test}}\}$. The classical training approach involves minimizing the following loss function,
$\mathcal{L}=\ell(f(\mathcal{G}_{n},\theta),y_{n}),$ (1)
where $\ell$ denotes the cross-entropy loss function.
To improve the generalization performance of GNNs, we introduce a graph data augmentation strategy. For each training graph $\mathcal{G}_{n}$ in the dataset, we generate $M$ augmented graphs, denoted as $\{\widetilde{\mathcal{G}}_{n,m},\widetilde{y}_{n,m}\mid m=1,\ldots,M\}$, where $M$ is the number of augmented graphs generated per training graph. These augmented graphs are obtained using a graph augmentation generator $A_{\lambda}$, parameterized by $\lambda$ as a mapping $A_{\lambda}:\mathcal{G}_{n},y_{n}\rightarrow A_{\lambda}\left(\mathcal{G}_{n},y_{n}\right)\in\mathbb{G}\times\mathbb{Y},$ where $\mathbb{G}$ denotes the space of all possible graphs, and $\mathbb{Y}$ is the label space. The generator $A_{\lambda}$ may be either deterministic or stochastic with a dependence on a prior distribution $\mathcal{P}(\lambda)$. Examples of such augmentation strategies can be found in Appendix H.
@@
We use the notation $\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}$ to represent an augmented graph sampled from the augmentation strategy $A_{\lambda}$, i.e., $(\widetilde{\mathcal{G}}_{n}^{m},\widetilde{y}_{n}^{m})\sim A_{\lambda}(\mathcal{G}_{n},y_{n})$. With the augmented data, the loss function is modified to account for multiple augmented versions of each graph,
@@
$\mathcal{L}^{\text{aug}}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\ell(f(\widetilde{\mathcal{G}}_{n}^{m},\theta),\widetilde{y}_{n}^{m})\right].$
@@
For simplicity, we denote the loss for the original graph by $\ell(f(\mathcal{G}_{n},\theta),y_{n})=\ell(\mathcal{G}_{n},\theta)$ and the loss for an augmented graph as,
@@
$\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\ell(f(\widetilde{\mathcal{G}}_{n}^{m},\theta),\widetilde{y}_{n}^{m})\right]=\ell^{\text{aug}}(\widetilde{\mathcal{G}}_{n},\theta).$
Via the law of large numbers, $\mathcal{L}^{\text{aug}}$ is empirically estimated,
$\mathcal{L}^{\text{aug}}$ $=\frac{1}{N}\sum_{n=1}^{N}\ell^{\text{aug}}(\widetilde{\mathcal{G}}_{n},\theta)$
$\simeq\frac{1}{NM}\sum_{n=1}^{N}\sum_{m=1}^{M}\ell(f(\widetilde{\mathcal{G}}_{n}^{m},\theta),\widetilde{y}_{n}^{m}).$
@@
To understand the impact of data augmentation on the graph classification performance, we analyze the effect of the augmentation strategy on the generalization risk $\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta)\right]$. More specifically, we want to study the generalization error,
$\eta=\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\text{aug}})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\star})\right],$
where $\theta_{\text{aug}}$ and $\theta_{\star}$ are the optimal GNN parameters for the augmented and non-augmented settings, respectively,
$\theta_{\star}=\operatorname*{arg\,min}_{\theta}\mathcal{L}_{\theta},~{}~{}~{}~{}\theta_{\text{aug}}=\operatorname*{arg\,min}_{\theta}\mathcal{L}^{\text{aug}}_{\theta},$
which can be estimated empirically as follows,
@@
$\hat{\theta}_{\text{aug}}$ $=\operatorname*{arg\,min}_{\theta}\frac{1}{NM}\sum_{n=1}^{N}\sum_{m=1}^{M}\ell(\widetilde{\mathcal{G}}_{n}^{m},\theta).$
By theoretically studying the generalization error $\eta$, we aim to quantify the effect of each augmentation strategy on the overall classification performance, providing insights into the benefits and potential trade-offs of data augmentation in graph-based learning tasks. In Theorem 3.1, we present a regret bound of the generalization error using Rademacher complexity defined as follows (Yin et al., 2019),
@@
where $\epsilon_{n}$ are independent Rademacher variables, taking values $+1$ or $-1$ with equal probability, $P_{\epsilon}$ is the Rademacher distribution, and $\Theta$ is the hypothesis class. Rademacher complexity is a fundamental concept in statistical learning, which indicates how well a learned function will perform on unseen data (Shalev-Shwartz & Ben-David, 2014). Intuitively, the Rademacher complexity measures the capacity of a GNN to fit random noise (Zhu et al., 2009), where a lower Rademacher complexity indicates better generalization.